# Revisiting Graph Neural Networks on Graph-level Tasks: Comprehensive Experiments, Analysis, and Improvements

### v1

@@ 4.5.1. Graph Classification Tasks
We evaluate baselines and our model on the graph classification tasks in Table 3. Firstly, node-based GNNs, such as GCN and GIN, cannot achieve satisfactory performance. These models learn node representations and use global pooling, which overlooks local graph structures, resulting in suboptimal outcomes compared to other GNN types. Secondly, pooling-based approaches propose to progressively reduce the graph size hierarchically and progressively learn the graph representations, which cannot achieve the best performance. This is because they simply keep the top-k most important nodes and concentrate similar nodes into a community to learn graph representation, ignoring local graph structure information. Thirdly, subgraph-based approaches effectively capture intricate local structures within the graphs, making them well-suited for bioinformatics applications where understanding detailed interactions is essential. However, their computational intensity and potential overfitting to subgraph patterns present notable challenges. Fourthly, graph learning-based approaches reconstruct graph structure and node features to remove noisy information, thereby enhancing performance on noisy datasets, such as social networks. Lastly, SSL-based approaches propose to learn generalizable representations based on pretraining on unlabeled graphs and fine-tuning on labeled graphs. Particularly, our proposed model achieves the best or comparable performance on all datasets. This is because we propose adaptive contrastive learning and sample subgraphs to learn local structure information, thereby learning a generalizable and informative graph representation.
@@ 4.5.2. Graph Regression Tasks
For graph regression, we mainly evaluate the ability of GNNs to capture the semantics and core patterns of graphs, based on cycle counting and path counting for each graph. Lower MAE and higher R2 values indicate better performance. As shown in Table 4, node-based (except GIN), pooling-based, GL-based, and SSL-based approaches (except our model) do not achieve satisfactory performance. This is because these approaches are not designed to enhance the expressivity of GNNs, i.e., they cannot distinguish whether two graphs are isomorphic or have the same cycles. In contrast, GIN, subgraph-based GNNs, and our model aim to theoretically increase the expressivity of GNNs, thereby achieving satisfying performance. Particularly, our model performs the best on all datasets. This is because our model is guaranteed to have better expressivity, such as counting 8-Cycle, while other approaches have low expressivity. Additionally, we propose a unified contrastive learning approach to augment graphs and achieve more generalizable graph representations.
4.6. Efficiency Evaluation
We evaluate the efficiency of GNNs in terms of time and GPU peak memory consumption during both the training and inference phases. As shown in Table 5, node-based GNNs are the most efficient because they directly learn node representations from the original graphs without any additional processing. Secondly, the pooling-based approach, TopK also achieves good efficiency and low memory usage by retaining only the most important nodes However, GMT and EPool are slightly less efficient as it computes use nodes features to compute the cluster representation and pairwise similarities between connected nodes to form clusters, respectively. Thirdly, subgraph-based and GL-based approaches are less efficient due to their reliance on additional graph processing. Subgraph-based GNNs sample subgraphs for each root node or generate multiple subgraphs and learn representations for each one, which increases time and memory costs. Similarly, GL-based approaches iteratively optimize the graph structure and node features while training the GNNs, further adding to resource consumption. Lastly, SSL-based approaches are computationally intensive in training because they augment graphs into positive pairs and optimize GNNs by maximizing the similarity between these pairs, which increases time and memory usage. However, during inference, they do not require graph augmentation and operate directly on the original graphs, resulting in efficiency comparable to node-based GNNs.
@@ 4.6. Efficiency Evaluation
Figure 3. Robustness evaluation.
@@ 4.7. More Scenarios
We evaluate GNNs in more realistic scenarios, including noisy graphs, imbalanced graphs, and few-shot graphs. To ensure clarity, we select one representative graph dataset from each domain for evaluation, i.e., IMDB-M, ENZYMES, NCI1, and 4-Cycle. Additionally, we compare the most effective GNN model from each GNN category, including GCN, GMT, with AK+ and MO for classification tasks, I2GNN and HGPSL for regression tasks, with our model.
4.7.1. Robustness Evaluation
We evaluate the robustness of existing GNNs on noisy and incomplete graphs. Specifically, for each graph $G(V,\mathbf{A},\mathbf{X})$, we randomly delete $\alpha||\mathbf{A}||_{0}$ existing edges by varying a ratio $\alpha\in\{0.1,0.2,0.3,0.4,0.5\}$. As shown in Figure 3, as $\alpha$ increases, the performance of GNNs decreases and our proposed model outperform other baselines under the same ratio $\alpha$. Specifically, node-based GCN shows the sharply decline due to their reliance on the original graph structure and global pooling, which fail to capture local patterns effectively. Also, Pooling-based approach GMT also performs unsatisfying as it separates nodes into different clusters based on node similarity and connectivity, while edge deletions significantly affect the cluster results. Subgraph-based approach AK+ performs better by capturing intricate local patterns, but their reliance on subgraph sampling makes them sensitive to edge deletions. GL-based method MO shows resilience by dynamically learning graph structures. Particularly, our proposed model outperforms all baselines, maintaining strong performance on all $\alpha$ values. By leveraging adaptive contrastive learning and subgraph sampling, it effectively captures local structures and mitigates the impact of noise, resulting in superior robustness.
@@ 4.7.2. Imbalance Data Evaluation
In the graph classification task, given a dataset $\mathcal{G}={(G_{i},y_{i})}$, we simulate class imbalance by setting the proportions of training samples per class as $\{1,\frac{1}{2^{\beta}},\frac{1}{3^{\beta}},\dots,\frac{1}{|\mathcal{Y}|^{\beta}}\}$, where $\beta\in\{0,0.5,1,1.5,2\}$ controls the imbalance ratio. The number of samples in the first class is fixed under all $\beta$ values. For graph regression tasks, where $y_{i}\in[y_{\text{min}},y_{\text{max}}]$, we partition the label range into three equal buckets and vary the proportions of the groups as $\{1,\frac{1}{2^{\beta}},\frac{1}{3^{\beta}}\}$ to simulate imbalance. A higher $\beta$ indicates a greater level of imbalance. As shown in Figure 4, GCN and GMT struggle with both classification and regression tasks due to their reliance on global graph aggregation, which dilutes the features of minority class nodes and edges in imbalanced datasets. Similarly, AK+ (and I2GNN) and MO (and HGP), despite incorporating subgraph-level information or reconstructing graph structures, fail to explicitly address class imbalance, limiting their ability to preserve minority class features under skewed distributions. In contrast, our model demonstrates superior robustness across all tasks. By effectively capturing task-relevant subgraphs and leveraging adaptive contrastive learning, it mitigates the dominance of majority class features and balances feature representations, preserving the features of minority classes even under severe imbalance.
@@ 4.7.3. Few-shot Evaluation
Specifically, For graph classification tasks, given a training graph dataset $\mathcal{G}=\{(G_{i},y_{i})\}$, we set the number of training graphs per class as $\gamma\in\{5,10,15,20,25\}$. For graph regression tasks, we partition the range $[y_{\text{min}},y_{\text{max}}]$ into five equal buckets to create discrete groups and sample $\gamma$ training instances for each bucket. As shown in Figure 5, GCN and GMT struggle in few-shot settings due to their reliance on global graph aggregation, which oversmooths features and fails to preserve task-relevant substructures. Similarly, AK+, I2GNN, MO, and HGP, while incorporating subgraph-level information, lack mechanisms to prioritize task-relevant features or adapt to low-data conditions, leading to suboptimal performance. In contrast, our model demonstrates superior performance across all tasks by explicitly capturing task-relevant subgraphs and leveraging adaptive contrastive learning, achieving robust generalization even with very limited training data.
5. Conclusion
In this paper, we systematically revisit existing GNNs for graph-level tasks, categorizing them into five types: node-based, hierarchical pooling-based, subgraph-based, graph learning-based, and self-supervised learning-based approaches. To address limitations in current evaluation practices, we proposed a unified evaluation framework that ensures comprehensive and fair comparisons by incorporating diverse datasets, multiple tasks, and challenging scenarios such as noisy, imbalanced, and few-shot graphs. Furthermore, we propose a novel GNN model that enhances expressivity through a $k$-path rooted subgraph approach and improves generalization via adaptive graph contrastive learning. Extensive experiments demonstrated the superior performance of our model across a wide range of graph-level tasks and twenty-seven datasets.

### v2

@@ 4.6.1. Graph Classification Tasks
As shown in Table 4, node-based GNNs, such as GCN and SAGE, cannot achieve satisfactory performance. These models learn node representations and use global pooling. Global pooling overlooks local structural information, which is critical for distinguishing graphs in bioinformatics (e.g., ENZYMES) and chemistry (e.g., MUTAG). PNA achieves comparatively good performance as it captures richer neighborhood distributions by leveraging multiple aggregators and degree-scalers. Secondly, pooling-based approaches, such as GMT and TopK, achieve competitive performance, particularly on social networks like COLLAB and REDDIT. These methods progressively reduce the graph size by grouping nodes or retaining only the most important ones, thereby preserving multi-level structural information. However, they are less effective on datasets where fine-grained local structures (e.g., motifs) play a critical role, such as ENZYMES and NCI1. Subgraph-based methods, such as ECS, AK+, and I2, demonstrate highly competitive performance on bioinformatics and chemistry datasets because they break graphs into meaningful substructures, enabling them to capture important patterns that other methods often miss. However, they are out-of-memory (OOM) on large graphs or high node counts, such as REDDIT and COLLAB. Fourthly, graph learning-based approaches, such as MO and HGP, perform well on noisy social datasets, such as IMDB-B and IMDB-M. By dynamically reconstructing graph structures and removing irrelevant edges or nodes, these approaches enhance robustness and improve generalization. However, they are less effective on datasets where the original graph structure is already well-formed molecular graphs, such as MUTAG. Lastly, SSL-based methods like MVG and GCA perform robustly across multiple datasets, leveraging pretraining on unlabeled graphs to learn generalizable representations. However, SSL-based approaches are computationally expensive due to graph augmentation and can lead to OOM issues.
@@ 4.6.2. Graph Regression Tasks
For graph regression tasks, the focus is primarily on evaluating how well GNNs can capture the semantics and key structural patterns of graphs, such as cycle and path counts for each graph. Lower MAE and higher $R^{2}$ scores reflect better performance. As shown in Table 5, node-based methods (excluding GIN and PNA), pooling-based models, GL-based techniques, and SSL-based approaches generally fail to deliver satisfactory results. This is because these methods are not specifically designed to enhance the expressiveness of GNNs, meaning they cannot effectively differentiate between isomorphic graphs or graphs with identical cycles. In contrast, GIN, PNA, and subgraph-based GNNs explicitly aim to improve the theoretical expressiveness of GNNs, resulting in superior performance. GIN and PNA aim to improve the theoretical expressiveness to approximate the Weisfeiler-Lehman (1-WL) isomorphism test. Subgraph-based models like ECS, AK+, and I2 consistently outperform other approaches on almost all regression datasets. By breaking graphs down into overlapping or rooted subgraphs, these methods capture rich structural details that enhance their theoretical expressivity. As a result, they are better able to distinguish between graph isomorphism classes and accurately identify important motifs, which leads to improved performance. Lastly, as the complexity of the target motif increases (e.g., from 3-Cycle to 8-Cycle), most GNNs (except subgraph-based ones) experience a drastic performance drop. This highlights the limitations of many GNN architectures in capturing higher-order dependencies and complex structural relationships.
@@
Figure 3. Robustness evaluation on noisy graphs.
@@
4.7. Efficiency Evaluation
We evaluate the efficiency of GNNs in terms of total training time, inference time, GPU peak memory usage during training and inference. As shown in Table 6, node-based GNNs exhibit the highest efficiency across all datasets. Their simplicity in aggregating neighbors’ information without additional graph processing ensures minimal time and memory usage. However, this efficiency comes at the cost of limited expressiveness as discussed in Section 4.6. Secondly, pooling-based methods strike a balance between efficiency and performance. TopK is efficient due to its node-pruning strategy, which reduces graph size while preserving key features. However, more advanced pooling approaches, such as GMT, are computationally expensive due to their clustering operations based on node similarity. Thirdly, subgraph-based GNNs (e.g., ECS, AK+, I2) are computationally intensive because they rely on generating multiple subgraphs for each graph. Despite their high resource requirements, these methods excel at capturing graph motifs and complex structural patterns, as shown in Section 4.6. Their computational cost makes them less efficient for real-time or large-scale applications. GL-based methods (e.g., VIB, HGP, and MO) dynamically reconstruct graph structures during training, which adds significant computational overhead. While they improve robustness to noise and graph quality, their iterative optimization process limits scalability to larger datasets. SSL-based methods (e.g., RGC, MVG, GCA) are computationally expensive during training because they require graph augmentations and additional contrastive learning stage.
In summary, node-based GNNs are the most efficient but lack expressiveness, while pooling-based models strike a balance between efficiency and performance. Subgraph-based and GL-based approaches offer superior expressiveness but suffer from high computational costs. SSL-based methods, though computationally expensive during training, are efficient during inference, making them suitable for pretraining scenarios.
@@ 4.8. More Scenarios
We further evaluate GNNs in three realistic and challenging scenarios, including noisy graphs, imbalanced graphs, and few-shot graphs. To ensure clarity, we select one representative graph dataset from each domain for evaluation, i.e., IMDB-M, ENZYMES, NCI1, and 4-Cycle. Additionally, we compare some of the most representative models from each GNN category, including GCN (node-based), GMT (pooling-based), and GCA (SSL-based). For top perform, we select AK+ (subgraph-based), MO (GL-based) for classification tasks, while I2 (subgraph-based) and HGP (GL-based) for regression tasks.
4.8.1. Robustness Evaluation on Noisy Graphs
[…]
@@ 4.8.2. Imbalance Data Evaluation
For the graph classification task, given a dataset $\mathcal{G}={(G_{i},y_{i})}$, we simulate class imbalance by adjusting the proportion of training samples in each class to follow the sequence $\{1,\frac{1}{2^{\beta}},\frac{1}{3^{\beta}},\dots,\frac{1}{|\mathcal{Y}|^{\beta}}\}$, where $\beta\in\{0,0.5,1,1.5,2\}$ determines the degree of imbalance. The sample count for the first class remains constant across all $\beta$ settings. For graph regression tasks, where $y_{i}$ spans the interval $y_{\text{min}},y_{\text{max}}]$, we divide the label range into three equal-sized buckets and set the group proportions as $\{1,\frac{1}{2^{\beta}},\frac{1}{3^{\beta}}\}$ to create imbalance, with larger $\beta$ corresponding to higher level of imbalance. As shown in Figure 4, accuracy drops across various GNN architectures as imbalance increases, particularly in datasets with minority classes. Node-based GNN (e.g., GCN) and pooling-based GNN (GMT) struggle because their global aggregation mechanisms tend to average out minority signals, making it difficult to distinguish rare patterns. Subgraph-based (AK+) and graph learning-based (MO) GNNs, while capturing richer structural information, do not explicitly address class imbalance, leading to performance degradation under extreme imbalance, especially for less classes. SSL-based model (GCA), although capable of leveraging abundant unlabeled data, likewise fails to correct imbalance on its own and converges to performance similar to other baselines unless augmented with imbalance-aware strategies (Liu et al., 2023).
@@ 4.8.3. Few-shot Evaluation
We simulate data scarcity by limiting the number of training samples. For classification tasks, we construct training sets where the number of labeled graphs per class is in $\gamma\in\{5,10,15,20,25\}$. For regression tasks, we partition the label into five equal-width buckets and uniformly sample $\gamma$ training instances per bucket. The results, as summarized in Figure 5, reveal that most GNN architectures exhibit significant performance degradation as $\gamma$ decreases. Node-based (GCN) and pooling-based (GMT) models are particularly affected, as their global aggregation mechanisms rely heavily on abundant labeled data to learn meaningful representations. Interestingly, while subgraph-based models such as AK+ and I2, as well as graph learning-based models like MO and HGP, are theoretically capable of leveraging local structural patterns, they do not demonstrate substantial resilience to data scarcity in practice. This is primarily because current implementations lack explicit mechanisms to identify and prioritize the most informative subgraphs or adaptively focus on critical features when labeled data is limited. Their performance improvements over standard baselines are thus marginal in few-shot scenarios, indicating that richer local modeling alone does not guarantee data efficiency (Wang et al., 2023c).
4.9. Correlation Analysis
4.9.1. Graph Topological Features
To investigate the intrinsic relationship between graph topology and GNN model performance, we characterize each graph using a comprehensive set of topological features. These features capture properties ranging from macroscopic global characteristics to fine-grained local patterns. We categorize them into three groups: (i) Global Structural Features, (ii) Local Structural Features, and (iii) Node Distributions.
Global Structural Features. These metrics provide a high-level summary of the entire graph’s connectivity and organization. They typically return a single scalar value for each graph.
  • Average Degree (Deg.) and Density (Den.). For a graph $G=(V,E)$, let $|V|$ and $|E|$ be the number of nodes and edges, respectively. $\text{Avg. Deg}=\frac{2|E|}{|V|}$. Density is the ratio of actual edges to possible edges:
    (16) $\text{Density}=\frac{2|E|}{|V|(|V|-1)}.$
    They measure the overall sparsity.
  • Average Shortest Path Length (SPL) and Diameter (Dia.). Let $d(u,v)$ denote the shortest path distance between nodes $u$ and $v$. $\text{SPL}=\frac{1}{|V|(|V|-1)}\sum_{u\neq v}d(u,v)$, and $\text{Diameter}=\max_{u\neq v}d(u,v)$. For disconnected graphs, these are computed on the largest connected component.
  • Degree Assortativity (Assr.). Calculated as the Pearson correlation coefficient between the degrees of connected nodes. Positive values indicate that high-degree nodes tend to connect with other high-degree nodes.
  • Modularity (Mod.). Quantifies the strength of community structure. A positive value indicates that connections within communities are denser than expected by random chance.
Local Structural Features. These metrics focus on small, recurring substructures. For node-level metrics, we aggregate them into graph-level features.
  • Average Clustering Coefficient (CC). For a node $v$, the local clustering coefficient $C_{v}$ is the density of the subgraph induced by its neighbors. We report the average over all nodes: $\text{Avg. CC}=\frac{1}{|V|}\sum_{v\in V}C_{v}$.
  • Triangle Counts (Tri.). Let $T_{v}$ be the number of triangles node $v$ participates in. Unlike global metrics, to capture both the prevalence and distribution of triangles, we report both the mean ($\mu$) and standard deviation ($\sigma$) of $T_{v}$ across all nodes. High $\mu$ implies a dense graph, while high $\sigma$ indicates a significant degree of heterogeneity.
Node Distributions. These features analyze the distribution of node importance. Since these are originally defined at the node level, we characterize the entire graph by reporting both the Mean ($\mu$) and Standard Deviation ($\sigma$) for each metric.
  • Betweenness Centrality (BC.). $C_{B}(v)$ represents the fraction of all-pairs shortest paths passing through node $v$. We report $\mu_{\text{BC}}$ and $\sigma_{\text{BC}}$. A high $\sigma_{\text{BC}}$ indicates the presence of critical ”bridge” nodes controlling information flow.
  • PageRank (PR.). Measures node influence based on transitive importance. We report $\mu_{\text{PR}}$ and $\sigma_{\text{PR}}$. A high $\sigma_{\text{PR}}$ reveals a hierarchical structure with distinct authority nodes.
  • K-Core Number (KC.). The core number $k_{v}$ is the largest $k$ such that $v$ belongs to a $k$-core. We report $\mu_{\text{KC}}$ and $\sigma_{\text{KC}}$. A high $\mu_{\text{KC}}$ indicates global robustness, while high $\sigma_{\text{KC}}$ reflects a core-periphery structure.
4.9.2. Evaluation
Given GNNs $\mathcal{M}=\{M_{i}\}_{i=1}^{n}$ and graph dataset $\mathcal{D}=\{G_{j}\}_{j=1}^{m}$, we construct a performance matrix $\mathbf{P}\in\mathbb{R}^{n\times m}$ where $\mathbf{P}[i][j]$ is the result of $M_{i}$ on $G_{j}$ (e.g., accuracy or MAE). We denote each $k$-th topological feature of $\mathcal{D}$ as $\mathbf{f}_{k}$, where $\mathbf{f}_{k}[j]$ is the value of the $k$-th feature for graph $G_{j}$. We assess the relationship between model performance and $k$-th feature with Spearman’s correlation $\rho(\mathbf{P}[i],\mathbf{f}_{k})$, with a significance threshold of $\alpha=0.05$.
4.9.3. Insights
Figure 6. Correlation between graph features and GNNs. Each cell shows the Spearman’s $\rho$ with * for $p<0.05$ and $\dagger$ for $0.05\leq p<0.1$.
As illustrated in Figure 6, we observe:
  • Structurally simpler models (e.g., Node-based) show weak correlations with most features, reflecting their reliance on local aggregation rather than global topology.
  • Graph density negatively correlates with most models, suggesting high connectivity exacerbates over-smoothing and noise.
  • High local clustering and assortativity negatively impact Subgraph-based and SSL-based models, likely due to the “rich-club” effect distracting from peripheral structures.
  • Conversely, graph sparsity and high Betweenness Centrality positively correlate with Hierarchical (e.g., HGP) and SSL-based models, demonstrating their ability to leverage clear structures for long-range information flow.
Most importantly, the absence of universally significant correlations confirms that (i) model selection cannot rely on a single structural indicator, and (ii) no single architecture consistently dominates. This underscores the critical importance of comprehensive benchmarks across diverse domains, as conducted in OpenGLT, to provide an empirical basis for scenario-specific model selection.
5. Conclusion and Future Direction
This paper presents a comprehensive experimental study on graph neural networks for graph-level tasks, categorizing existing models, introducing a unified evaluation framework (OpenGLT), and benchmarking 16 GNNs across 26 diverse datasets under challenging scenarios such as noise, imbalance, and limited data. Our results reveal that no single model excels universally, highlighting important trade-offs between expressiveness and efficiency, and emphasizing the need for robust evaluation in realistic conditions.
Based on these findings, promising future directions include the development of scenario-adaptive or hybrid GNN architectures that dynamically leverage different model strengths, research into lightweight and scalable algorithms for practical deployment, and the incorporation of transfer and foundation model techniques to improve generalization and data efficiency, particularly when labeled data is scarce. Our study provides valuable benchmarks and guidance for advancing GNN research on graph-level tasks.